find the molarity of 3m solution of NaCl having density 1.05 gm/ml.
 

Grade:11

1 Answers

Seema tiwari
5 years ago
Molarity = 3 mol L^-1 
Mass of NaCl = 1000 × density = 1000 × 1.25 = 1250 g
Mass of water in solution = 1250 - 175.5 g
Hence , molarity (m) = moles of solute / mass of solvent in kg = 3/1.0745 = 2.79m